OpenStudy (anonymous):

Please help me with 1. what is the solution of the system? solve using matrices { 7x+2y=-38 {13x-3=32 A. [-54] [-10] B. no solution C. [10] [54] D. [-10] [-54]

OpenStudy (anonymous):

@hartnn @GamerGirl.1 @grace111 @ganeshie8 @Luigi0210 @hoblos @hello1213 @lyubas @surjithayer @skullpatrol

OpenStudy (anonymous):

im sorry I cant figure this out :(

OpenStudy (anonymous):

its find ill try and figure it out :/

OpenStudy (anonymous):

Do you want to do it using a formula (involving finding the inverse of a matrix) or by reducing an augmented matrix?

OpenStudy (anonymous):

yah i do i just dont understand how to do it

OpenStudy (anonymous):

We'll do it by row reducing. Did you forget a y in the second equation?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think i may have figured it out is it [10] [54]

OpenStudy (anonymous):

That does not satisfy the second equation: \(13x-3y = \begin{cases}13(10)-3(54) = 130 - 112 = 18 \\ 32\end{cases}\)

OpenStudy (anonymous):

Sorry, it should be \[13x-3y = \begin{cases}13(10)-3(54) = 130 - 162 = -32 \\ 32\end{cases}\] Did you forget a -?

OpenStudy (anonymous):

oh i mixed it up

OpenStudy (anonymous):

I guess you are right then. What method did you use?